# A machine-learning classifier for the postmerger remnant of binary neutron stars

### v1

@@ Acknowledgements.
We thank Chris van den Broeck, Brian Healy, Thibeau Wouters, Vsevolod Nedora, and Sebastian Kahn for the helpful discussion. The authors acknowledge funding from the Daimler and Benz Foundation for the project “NUMANJI” and from the European Union (ERC, SMArt, 101076369). Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation, as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. KAGRA is supported by Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) in Japan; National Research Foundation (NRF) and Ministry of Science and ICT (MSIT) in Korea; Academia Sinica (AS) and National Science and Technology Council (NSTC) in Taiwan.
@@ Appendix A Assessing the probability of misclassification
    where $\left\{\mathbf{x}\right\}$ is the set of inputs, $K$ the total number of classes, and $y=\left\{\omega_{1},...,\omega_{K}\right\}$ the labels.
@@
In the following, we show how we can derive the knowledge uncertainty with the Bayesian ensemble method, considering as an example Classifier A. We generate an ensemble of 100 models and compute the knowledge uncertainty on each point in the validation set as described in Eq. 14. The underlying idea is that larger (smaller) values of $\mathcal{U}(\mathbf{x})$ indicate that the prediction for the point $\mathbf{x}$ is more (less) likely to be a misclassification. Therefore, following Ref. [101], we can use the uncertainty $\mathcal{U}(\mathbf{x})$ to consider a point $\mathbf{x}$ correctly classified (label 0) or, instead, misclassified (label 1), based on an threshold $T$:
@@
Looking at our validation set, for which we know if the points are correctly classified or not, we check if, for different thresholds $T$, the criterion in Eq. 15 would have correctly identified them as misclassified or not, and we compute the area under the Receiver-Operating-Characteristic curve (AUROC) [104]. Since ROC curves include only information about true and false positive rates, the AUROC values can be biased towards 1.0 if the dataset includes significantly more negatives than positives. Therefore, we also calculate the value for area under the precision-recall curve (AUPR), where the recall is the true positive rate and the precision the fraction of true positives compared to the total number of points labeled as positives [105, 106]. For both AUROC and AUPR, the closer their value is to 1, the more robust the classifier is. For the example of Classifier A with an ensemble of 100 models, we find AUROC = 0.9875 and AUPR = 0.79.
In the event of a detection, therefore, we can evaluate whether the related prediction is likely a misclassification, and therefore more investigations are needed, by looking at its uncertainty $\mathcal{U}(x)$ in relation to an optimal threshold value. However, considering every time an ensemble of models is computationally expensive and unnecessarily elaborate given that, as already mentioned, NR errors dominate our uncertainties anyway. Therefore, we perform a similar check, but considering only one model, i.e., Classifier A, not an ensemble. We compute the data uncertainty on this single model as in Eq. 13, and we build an ROC curve following the same procedure described above. We look for the optimal threshold value $T_{\rm\textsc{a}}$, which keeps the false positives rate below 0.1 and the true positive rate close to 1, and we find $T_{\rm\textsc{a}}\sim 0.3$. With the same method, we look for the ideal threshold for Classifier B, $T_{\rm\textsc{b}}$. In this case, if we want to keep the false positive rate small ($\sim 0.1$), we need an uncertainty threshold 0.67, which corresponds to just above $50\%$ true positive rates; if, instead, we want a true positive rate close to $100\%$, we need to set the uncertainty threshold to 0.06, which, however, implies a quite high false positive rate (however, such a conservative choice is acceptable since the goal of this uncertainty threshold is to identify inputs that would require more study). Finally, repeating the procedure for Classifier C, we conclude that setting a threshold $T_{\rm\textsc{c}}=0.09$ allows us to have the true positive rate basically at 1, while keeping the false positive rate $\sim 0.2$.

### v2

@@ Acknowledgements.
We thank Chris Van Den Broeck, Brian Healy, Thibeau Wouters, and Vsevolod Nedora for the helpful discussion. The authors would like to thank Sebastian Khan for suggesting the use of interpretable machine learning techniques, such as SHAP, to enhance the understanding of our model and analysis. The authors acknowledge funding from the Daimler and Benz Foundation for the project “NUMANJI” and from the European Union (ERC, SMArt, 101076369). Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation, as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. KAGRA is supported by Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) in Japan; National Research Foundation (NRF) and Ministry of Science and ICT (MSIT) in Korea; Academia Sinica (AS) and National Science and Technology Council (NSTC) in Taiwan.
@@ Appendix A Assessing the probability of misclassification
    where $\left\{\mathbf{x}\right\}$ is the set of inputs, $K$ is the total number of classes, and $y=\left\{\omega_{1},...,\omega_{K}\right\}$ are the labels.
@@
In the following, we show how we can derive the knowledge uncertainty with the Bayesian ensemble method, considering as an example Classifier A. We generate an ensemble of 100 models and compute the knowledge uncertainty $\mathcal{U}(\mathbf{x})$ on each point $\mathbf{x}$ in the validation set as described in Eq. 14. The underlying idea is that larger (smaller) values of $\mathcal{U}(\mathbf{x})$ indicate that the prediction for the point $\mathbf{x}$ is more (less) likely to be a misclassification. Therefore, following Ref. [112], we can use the uncertainty $\mathcal{U}(\mathbf{x})$ to consider a point $\mathbf{x}$ correctly classified (label 0) or, instead, misclassified (label 1), based on a threshold $T$:
@@
Looking at our validation set, for which we know if the points are correctly classified or not, we check if, for different thresholds $T$, the criterion in Eq. 15 would have correctly identified them as misclassified or not, and we compute the area under the Receiver-Operating-Characteristic curve (AUROC) [116]. Since ROC curves include only information about true and false positive rates, the AUROC values can be biased toward 1.0 if the dataset includes significantly more negatives than positives. Therefore, we also calculate the value for the area under the precision-recall curve (AUPR), where the recall is the true positive rate and the precision is the fraction of true positives compared to the total number of points labeled as positives [117, 118]. For both AUROC and AUPR, the closer their value is to 1, the more robust the classifier is. For the example of Classifier A with an ensemble of 100 models, we find AUROC = 0.9875 and AUPR = 0.79.
In the event of a detection, therefore, we can evaluate whether the related prediction is likely a misclassification, and therefore more investigations are needed, by looking at its uncertainty $\mathcal{U}(x)$ in relation to an optimal threshold value. However, considering an ensemble of models every time is computationally expensive and unnecessarily elaborate given that, as already mentioned, NR errors dominate our uncertainties anyway. Therefore, we perform a similar check, but considering only one model, i.e., Classifier A, not an ensemble. We compute the data uncertainty on this single model as in Eq. 13, and we build an ROC curve following the same procedure described above. We look for the optimal threshold value $T_{\rm\textsc{a}}$, which keeps the false positive rate below 0.1 and the true positive rate close to 1, and we find $T_{\rm\textsc{a}}\sim 0.3$. With the same method, we look for the ideal threshold for Classifier B, $T_{\rm\textsc{b}}$. In this case, if we want to keep the false positive rate small ($\sim 0.1$), we need an uncertainty threshold 0.67, which corresponds to just above a $50\%$ true positive rate; if, instead, we want a true positive rate close to $100\%$, we need to set the uncertainty threshold to 0.06, which, however, implies a quite high false positive rate (however, such a conservative choice is acceptable since the goal of this uncertainty threshold is to identify inputs that would require more study). Finally, repeating the procedure for Classifier C, we conclude that setting a threshold $T_{\rm\textsc{c}}=0.09$ allows us to have the true positive rate basically at 1, while keeping the false positive rate $\sim 0.2$.